The perimeter of a rectangle is 524 in. The length is 16cm greater than the width. What is the width?
Answer by Alan3354(30993)

L + W = 524/2 = 262
------
262/2= 131
16/2 = 8
--> 131 - 8 & 131 + 8
W = 123
L = 139